OpenStudy (aduller):

Find the 62nd term of the following arithmetic sequence 3,7,11,15,....

OpenStudy (21younga):

You have to add 4 to 3 62 times and that's how you get your answer.

OpenStudy (mathmale):

A formula applying to arithmetic sequences that I learned decades ago is l=a+(n-1)d, where l is the last term, a is the first term, n is the number of terms and d is the difference from one term to the next following term.

OpenStudy (mathmale):

Find the 62nd term of the following arithmetic sequence 3,7,11,15,.... Let's check whether this formula "works." In the arith sequence given above, a=3, l=7 and d=4. We should be able to predict that the 4th term is 15. l=a + (n-1)d becomes l=3+(4-1)4 = 3+(3)4=15. correct. Now predict the 62nd term. Note: Please let @aduller have first stab at this!